You have a summer job that pays time and a half for overtime. That is, if you work more than 40 hours a week, your hourly wage for the extra hours is 1.5 times your normal hourly wage of $7.
a. Write and graph a piecewise function that gives your weekly pay "P" in terms of the number "H" of hours you work.
b. How much will you get paid if you work 45 hours?

Answer by addingup(3677)

a.
P = 7(40)+(7*1.5)(H-40)
-----
b. - for 45 hrs
P = 7(40)+(7*1.5)(45-40)
P = 7(40)+(10.5*5)
Use your calculator to get your answer. First multiply on both sides of the + sign, then add and you are done.